Fermi National Laboratory


Accelerator Update

Wednesday March 31
The day shift began with Operations monitoring store 3335, with the Antiproton Source (Pbar) stacking, with vacuum technicians accessing Switchyard (SY) F1 area, and with MiniBooNE taking beam.

At 9:27 AM, techs and experts accessed the Main Injector (MI) to look for and repair LCW leaks. Recycler experts also accessed MI to inspect a correction element. And at the same time techs accessed the Booster to investigate beam-monitoring problems.

Pbar Debuncher studies began at 10:28 AM.

At 11:19 AM, a water expert reported finding and temporarily repairing a LCW leak on a quadrupole in the P1 transfer line between MI and the Tevatron (TeV).

Booster techs came out of access at 12:35 PM.

Stacking resumed at 1:44 PM.

Techs in SY F1 area reported at 2:04 PM that they had replaced a SWIC. They continued to work on the Septa vacuum.

A MI expert began a MECAR study at 2:41 PM. Stacking stopped during this study.

A TeV expert began end-of-store separator scans at 3:22 PM.

The MI MECAR study ended and stacking resumed at 5:03 PM.

The TeV separator study ended and a TeV collimator study began at 5:26 PM.

Vacuum techs in SY F1 completed their work at 5:48 PM. An expert began conditioning the Septa. He then established beam to the SY dump at 6:22 PM.

At 6:31 PM, a Pbar Debuncher RF station, (DRF1) tripped off. Operators investigated and discovered a blown fuse. They replaced the entire power supply. Unfortunately the fan for this supply wasn't working. The supply was replaced again later.

TeV experts completed their collimator study at 7:28 PM

Operations terminated store 3335 at 7:32 PM.

Operations put the TeV through a dry squeeze at 7:44 PM.

Operations put the TeV into shot setup at 8:42 PM and at 10:40 PM established store 3337 with an initial luminosity of 55.10E30. Pbar resumed stacking.

At 10:59 PM, Operations began trying to resolve TeV colliding beam problems.

At 11:55 PM, Operations discovered a TeV collimator that hadn't moved with the others.

Thursday April 1
The midnight shift began with Operations monitoring store 3337, with Pbar stacking, with operators conducting SY studies, with beam to MiniBooNE, and with the TeV collimator move in progress.

At 12:41 AM, Operations reported that the TeV collimator move helped resolve enough of the colliding beam problems that the Collider Detector at Fermilab (CDF) was able to turn its detector back on.

The Pbar Debuncher bunch rotation stopped operating properly at 6:06 AM. Operators investigated and then called experts. At 6:52 AM, and expert found a loose connection and bunch rotation resumed normal operation.

Operations established beam to the Meson MTest experiment at 8:53 AM.

At 9:08 AM, a Linac RF station (LRF5) tripped off and wouldn't reset. Techs replaced a contactor coil and had the station back on line about an hour later.

Pbar resumed stacking at 10:25 AM.

A Booster expert began a collimator study at 12:45 PM.

Operations shut down all SY and Meson beam at 1:13 PM for access and vacuum work.

At 2:47 PM, the TeV suffered a six-house quench at sectors A1, A4, BA, BB, BD, and F4. The sector A0 antiproton abort kicker prefired and caused the quench.

Operations prepared the CDF and D-zero (D0) collision halls for access at 3:30 PM.

The vacuum techs completed their access into SY F1 at 4:32 PM.

Both CDF and D0 were out of access by 7 PM.

Operations began establishing beam to MCenter at 7:43 PM.

Operations put the TeV through a wet squeeze at 9:06 PM.

Operations put the TeV into shot setup at 10:57 PM.

Friday April 2
The midnight shift began with the TeV in shot setup, with SY off until the new store is established, and with beam to MiniBooNE.

At 12:03 AM, a TeV RF station tripped off due to its upstream water pump tripping off.

At 12:13 AM, Operations established store 3341 with an initial luminosity of 40.76E30. Pbar resumed stacking. SY resumed taking beam.

At 6:03 AM, the TeV suffered a multi-house quench. Operations and TeV experts investigated the cause of the quench.

Today's Plans
The plans for today include first finding the cause for the multi-house quench and then to allow the TeV to conduct orbit studies. Operations will establish a new store after all studies have been completed.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/2/2004   email Fermilab